OpenStudy (anonymous):

aye can sum1 help me

OpenStudy (immanuelv):

with math?

OpenStudy (anonymous):

ya

OpenStudy (immanuelv):

I can try depending on the question

OpenStudy (anonymous):

Write an exponential function y = abx for a graph that includes (–4, 72) and (–2, 18). f(x) = 2(4)x f(x) = 0.5(4.5)x f(x) = 4.5(0.5)x f(x) = 4(0.2)x

OpenStudy (immanuelv):

do you have a calculator? ;)

OpenStudy (immanuelv):

graphing calculator

OpenStudy (anonymous):

nah

OpenStudy (immanuelv):

what you need to do is plug in those coordinates for instance you were given (-4, 72) so plug in -4 into x and 72 into f(x) for each function... thats a start

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Good idea to use calculator, but if wana show your work then solving it is completely different